# DRWKV: Focusing on Object Edges for Low-Light Image Enhancement

### v2

@@ III-D Loss Function
RetinexNet [21] 587.47 15.61 0.414 7.235 16.10 0.407 9.425 17.14 0.756 5.405 20.84 0.617 7.924 20.96 0.629 9.947
KinD [32] 34.99 15.77 0.548 5.517 18.67 0.772 9.221 15.22 0.542 6.124 20.98 0.597 7.221 21.65 0.621 8.714
RRDNet [33] 2.10 14.66 0.541 7.725 15.21 0.514 9.912 16.67 0.667 4.894 21.07 0.601 6.921 20.04 0.647 8.997
Zero-DCE [8] 5.21 14.86 0.559 4.215 14.12 0.512 8.652 14.93 0.531 5.507 21.23 0.752 6.941 20.39 0.691 8.014
SCI [34] 0.06 14.78 0.526 3.667 17.30 0.541 8.077 14.96 0.721 4.899 20.43 0.714 6.574 21.33 0.701 7.778
LLFormer [35] 22.52 16.23 0.642 3.947 20.56 0.801 9.162 24.42 0.914 4.614 25.66 0.832 6.627 28.45 0.821 8.011
UHDFormer [36] 48.37 21.07 0.604 4.621 21.59 0.804 7.230 22.60 0.903 5.476 28.13 0.875 7.021 22.75 0.732 6.657
Retinexformer [31] 15.51 22.24 0.701 2.976 21.65 0.835 4.735 25.10 0.925 3.971 28.96 0.879 5.441 28.96 0.896 7.201
IAT [37] 1.44 20.12 0.694 4.217 20.30 0.752 5.232 22.96 0.856 5.512 19.97 0.713 4.177 19.97 0.711 6.417
EnlightenGAN [38] 61.01 17.42 0.651 3.176 18.35 0.617 6.591 17.21 0.725 4.631 20.10 0.602 4.596 19.67 0.623 5.147
MambaIR [39] 60.66 21.14 0.704 3.004 20.11 0.802 4.928 24.75 0.922 4.522 25.11 0.873 3.661 26.35 0.510 4.477
RetinexMamba [40] 34.75 20.88 0.629 3.104 22.34 0.826 4.771 24.71 0.932 4.503 28.21 0.893 3.976 26.22 0.866 4.011
MambaLLIE [41] 20.85 20.64 0.627 3.047 22.14 0.821 4.021 24.81 0.940 4.071 29.74 0.902 4.417 28.92 0.869 4.417
DRWKV (Ours) 1.67 22.34 0.703 2.944 24.12 0.832 3.926 25.02 0.947 3.941 30.26 0.922 3.441 28.96 0.891 3.954
The newly defined $MS^{2}\text{-Loss}$ focuses on five key attributes of the model: structure, edge, illumination, artifact, and weight.
Decompositional consistency loss: Inspired by compressed sensing theory [42], it applies the L1 norm to image reconstruction:
$L_{\text{recon}}=\ |I-\hat{I}\|_{1},$ (9)
where $\|\cdot\|_{p}$ denotes the $p$-norm. $L_{\text{recon}}$ focuses on balancing pixel-level errors and structural information.
@@
$L_{\text{smooth}}=|\nabla L|\cdot\exp(-\lambda|\nabla I|),$ (11)
@@
$L_{\text{artifact}}=\|S\|_{1}+\delta\cdot\text{TV}(S),$ (12)
@@
$L_{\text{reg}}=\alpha^{2}+\beta^{2}+\gamma^{2}$ (13)
@@
$\begin{split}\mathcal{L}_{\text{MS2Loss}}&=\lambda_{1}L_{\text{recon}}+\lambda_{2}L_{\text{sparse}}+\lambda_{3}L_{\text{smooth}}\\ &\quad+\lambda_{4}L_{\text{artifact}}+\lambda_{5}L_{\text{reg}},\end{split}$ (14)
@@
IV Experiments
IV-A Experimental Settings
Implementation details: We implement model training and inference on the NVIDIA A800-80GB platform. Specific parameter settings are as follows: the number of training epochs $E$ is set to 500; the input image size $S$ is 128×128; the block numbers of the DRWKV structure are set to $N_{1}=4$ and $N_{2}=8$ respectively, and the input channel number $C$ is 16. The model optimizer uses Adam [43] with parameters set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$; the initial learning rate is $2\times 10^{-4}$, which is gradually reduced to $1\times 10^{-6}$ via the cosine annealing strategy [44].
Datasets: (1) LSRW-Huawei is a subset of the LSRW [45] dataset consisting of images captured by Huawei smartphones, which includes 2450 pairs of training images and 30 pairs of test images. (2) LOLv2 [18] is divided into two subsets: LOLv2 Real and LOLv2 Synthetic. The former is a dataset of real low-light scenes, containing 689 pairs of training images and 100 pairs of test images, covering various environments such as campuses and clubs. The latter is a synthetic dataset, consisting of 900 pairs of training images and 100 pairs of test images. (3) SDSD [46] dataset has two subsets: indoor and outdoor. The indoor subset includes 62 pairs of low-light/normal-light videos for training and 6 pairs of low-light/normal-light videos for testing. The outdoor subset contains 116 pairs of low-light/normal-light videos for training and 10 pairs of low-light/normal-light videos for testing.
Metrics: We use traditional quality (distortion) metrics PSNR and SSIM [47] to evaluate the results of low-light image enhancement, aiming to measure the pixel-level similarity and distortion between the enhanced images and reference images. Additionally, NIQE [48] is employed to assess the no-reference perceptual quality of the enhanced images.
IV-B Comparative Experiments
@@
Based on the quantitative results across five test benchmarks, the DRWKV model achieves performance improvements over multiple baseline methods with extremely low GFLOPS, as shown in Table I. Specifically, DRWKV exhibits superior performance in terms of naturalness on LSRW–Huawei, characterized by low baseline values of NIQE and a high optimization rate (1.1% higher than the second-place method). Additionally, on LOLv2, DRWKV shows differences in optimization strengths between the Real and Synthetic test benchmarks: it achieves the optimal model performance with an average improvement of 5.00 dB in PSNR on Real and 0.152 in SSIM on Synthetic, balancing the enhancement of image quality and the preservation of structural information. In low-light scenario videos, by independently enhancing low-light frames in SDSD, the model demonstrates average PSNR differences of up to 9.42 dB/8 dB in indoor and outdoor scenes, ensuring high ”restoration accuracy” of image details. For visualization results, refer to Figure 3; the generated results of our DRWKV model exhibit continuous object edges and high-fidelity details, which are highly consistent with the subject-focused edges.
TABLE II: Comparison of different Block combinations on the LOLv2 Real benchmark dataset.
@@
Retinex Configuration Results: This experiment focuses on the contradiction of edge distortion under low light. In terms of network structure, replaceable structures Block1 and Block2 are preset to assist in configuring Retinex, enabling DRWKV to achieve global optimum. Since the input end of the GER theory needs to receive the edge feature $E$ (a parameter from the output end of DRWKV), Block1 is constrained by structural limitations. To minimize disturbance to the network architecture, one-sided fixation is ultimately adopted. Deploying the GER theory into Block2 results in remarkable improvement rates of 6.9%/12.3%/4.8% for PSNR, SSIM, and NIQE, respectively, as shown in Table II. This strongly validates the core idea of gradient-guided global-edge coupling in the GER theory, demonstrating its strong practicality in addressing edge issues of low-light images and further confirming Hypothesis(1).
IV-C Ablation Studies
Loss Performance Results: A targeted ablation study was performed on the proposed loss function combinations. In Figure 4(a), the absence of the decompositional consistency loss $L_{recon}$ results in an overall darker reconstructed image with incomplete edge information, indirectly validating its role in structural and luminance reconstruction. In Figure 4(b), removing the edge sparsity loss $L_{sparse}$ leads to false edges generated by Gaussian noise contaminating flat regions, highlighting its contribution to preserving and enhancing low-light edges. Figures 4(c) and 4(d) illustrate uneven illumination and over-denoising effects caused by the removal of the illumination smoothness loss $L_{smooth}$ and artifact suppression loss $L_{artifact}$, respectively, confirming their balancing roles. Figure 4(e) demonstrates model instability with negative image artifacts when the parameter regularization loss $L_{reg}$ is removed, indicating that effective regularization can improve the model’s adaptability to unseen data. Additionally, the loss parameters for this model are specified as follows: $\lambda_{1}=1$, $\lambda_{2}=0.01$, $\lambda_{3}=0.1$, $\lambda_{4}=0.05$, and $\lambda_{5}=1e-4$.
@@ IV-C Ablation Studies
Component Ablation Results: This experiment aims to verify the effectiveness of the Evolving Scanning technique proposed in Hypothesis(2) for the edges of low-light images. Structurally, a progressive component integration strategy is adopted, with LOLv2 Real as the test benchmark and the basic structure as the test framework, to construct a five-level progressive validation system. The component configurations, ablation variables, and validation results at each level are shown in Table III. In terms of the basic structure, we retain the ViT-style block stacking structure of VRWKV and introduce the SSIM+L1 loss, enabling it to initially possess the capability to handle low-light issues. However, the model exhibits nearly ineffective feature extraction in illumination-degraded scenarios, characterized by extremely low SSIM (0.415) and PSNR (12.57 dB) values.
@@
Validation (3): Adding a shallow illumination restoration stage leads to a slight increase in image quality evaluation metrics, demonstrating that this stage contributes to improving pixel errors and structures.
@@
Validation (5): Replacing the SSIM+L1 loss with the proposed $MS^{2}\text{-Loss}$ specifically optimizes the training process, with SSIM and PSNR optimized to 0.832 and 24.12 dB, respectively.
@@
TABLE III: Ablation study on the LOLv2 Real benchmark dataset. If the RWKV blocks and training losses are not selected, Bi-RWKV and SSIM+L1 loss will be used by default. ✔
denotes the use of a ViT-style block stacking structure, whereas ✔ indicates the adoption of a UNet-like architecture.
ES-RWKV SIR Bi-SAB MS2Loss SSIM↑ PSNR↑ Params.(M) GFLOPs
@@
IV-D Extended Validation
To evaluate generalization, we applied DRWKV to low-light multi-object tracking on the UAVDark135 benchmark [49], comprising 135 sequences across diverse scenarios at 30 FPS and 1920×1080 resolution.
@@ IV-D Extended Validation
TABLE IV: Extended validation results of DRWKV on low-light object tracking performance. MOTA [50], IDF1 [51] and HOTA [52] are common performance evaluation metrics for object tracking.
@@
ByteTrack [53] 23.434 62.776 50.117 28.623 65.852 52.822
BoT-SORT [54] 18.242 62.916 49.556 25.496 64.049 51.435
Hybrid-SORT [55] 20.684 62.054 49.196 29.355 66.365 52.905
OC-SORT [56] 21.062 62.967 49.469 28.102 66.000 52.501
Deep OC-SORT [57] 19.514 63.205 49.857 24.929 58.016 49.124
@@
V Conclusion
In this study, we introduced the DRWKV model to address the challenge of preserving edge details in low-light image enhancement. By integrating the Global Edge Retinex (GER) theory, the model decouples illumination and edge structures, ensuring high-fidelity edge continuity even in extreme low-light conditions. The Evolving WKV Attention mechanism enhances spatial edge continuity, while the Bilateral Spectrum Aligner (Bi-SAB) and MS²-Loss optimize luminance and chrominance alignment, reducing artifacts and improving visual quality. Extensive experiments on benchmark datasets show that DRWKV outperforms existing methods in PSNR, SSIM, and NIQE, with lower computational cost. Additionally, it proves effective in tasks like low-light multi-object tracking, enhancing performance, especially in identity preservation. Future work will focus on adapting DRWKV for mobile platforms and further improving its robustness across various low-light scenarios, bridging the gap between academic research and real-world deployment.
@@ VI Supplementary Experiments
VII Architecture Details of DRWKV
@@ VII-A Light Preprocessing
Figure 5: Evolving Scanning RWKV’s structure, it consists of two parts: Evolving Spatial Mix and Channel Mix.
VII-B Evolving Scanning RWKV
Unlike traditional spatial modelling approaches, the proposed ES-RWKV breaks through spatial-size constraints, preserves structural integrity, enhances modelling capacity, and enables cross-space information interaction while markedly reducing the computational overhead of handling complex spatial features, thereby offering an efficient solution for spatial feature extraction and fusion in vision tasks. The ES-RWKV block is decomposed into two sequential stages: Evolving Spatial Mix and Channel Mix. The Evolving Scanning RWKV structure is shown in Figure.5
VII-B1 Evolving Spatial Mix

### v3

@@ 3.4 Loss Function
RetinexNet [wei2018deep] 0.8 587.47 15.61 0.414 7.235 16.10 0.407 9.425 17.14 0.756 5.405 20.84 0.617 7.924 20.96 0.629 9.947
KinD [zhang2019kindling] 8.0 34.99 15.77 0.548 5.517 18.67 0.772 9.221 15.22 0.542 6.124 20.98 0.597 7.221 21.65 0.621 8.714
RRDNet [9102962] 0.1 2.10 14.66 0.541 7.725 15.21 0.514 9.912 16.67 0.667 4.894 21.07 0.601 6.921 20.04 0.647 8.997
Zero-DCE [guo2020zero] 0.7 5.21 14.86 0.559 4.215 14.12 0.512 8.652 14.93 0.531 5.507 21.23 0.752 6.941 20.39 0.691 8.014
SCI [ma2022toward] 0.09 0.06 14.78 0.526 3.667 17.30 0.541 8.077 14.96 0.721 4.899 20.43 0.714 6.574 21.33 0.701 7.778
LLFormer [wang2023ultra] 13.2 22.52 16.23 0.642 3.947 20.56 0.801 9.162 24.42 0.914 4.614 25.66 0.832 6.627 28.45 0.821 8.011
UHDFormer [wang2024correlation] 0.3 48.37 21.07 0.604 4.621 21.59 0.804 7.230 22.60 0.903 5.476 28.13 0.875 7.021 22.75 0.732 6.657
Retinexformer [cai2023retinexformer] 1.6 15.51 22.24 0.701 2.976 21.65 0.835 4.735 25.10 0.925 3.971 28.96 0.879 5.441 28.96 0.896 7.201
IAT [cui2022you] 0.09 1.44 20.12 0.694 4.217 20.30 0.752 5.232 22.96 0.856 5.512 19.97 0.713 4.177 19.97 0.711 6.417
WalMaFa [Tan_2024_ACCV] 39.7 14.41 21.04 0.698 3.112 22.49 0.851 4.365 25.10 0.945 4.275 29.67 0.915 3.541 28.94 0.891 3.955
MambaIR [guo2024mambair] 20.4 60.66 21.14 0.704 3.004 20.11 0.802 4.928 24.75 0.922 4.522 25.11 0.873 3.661 26.35 0.510 4.477
RetinexMamba [bai2024retinexmamba] 4.6 34.75 20.88 0.629 3.104 22.34 0.826 4.771 24.71 0.932 4.503 28.21 0.893 3.976 26.22 0.866 4.011
MambaLLIE [weng2024mamballie] 4.4 20.85 20.64 0.627 3.047 22.14 0.821 4.021 24.81 0.940 4.071 29.74 0.902 4.417 28.92 0.869 4.417
DRWKV (Ours) 9.7 1.67 22.34 0.703 2.944 24.12 0.832 3.926 25.02 0.947 3.941 30.26 0.922 3.441 28.96 0.891 3.954
The newly defined MS2-Loss focuses on five key attributes of the model: structure, edge, illumination, artifact, and weight.
Decompositional consistency loss: Inspired by compressed sensing theory [doi:10.1126/science.1127647], it applies the L1 norm to image reconstruction:
$\mathcal{L_{\text{recon}}}=\|I-\hat{I}\|_{1},$ (11)
where $\|\cdot\|_{p}$ denotes the $p$-norm. $\mathcal{L_{\text{recon}}}$ focuses on balancing pixel-level errors and structural information.
@@
$\mathcal{L_{\text{smooth}}}=|\nabla L|\cdot\exp(-\lambda|\nabla I|),$ (13)
@@
$\mathcal{L_{\text{artifact}}}=\|S\|_{1}+\delta\cdot\text{TV}(S),$ (14)
@@
$\mathcal{L_{\text{reg}}}=\alpha^{2}+\beta^{2}+\gamma^{2}.$ (15)
@@
$\begin{split}\mathcal{L}_{\text{MS${}^{2}$-Loss}}&=\lambda_{1}\mathcal{L_{\text{recon}}}+\lambda_{2}\mathcal{L_{\text{sparse}}}+\lambda_{3}\mathcal{L_{\text{smooth}}}\\ &\quad+\lambda_{4}\mathcal{L_{\text{artifact}}}+\lambda_{5}\mathcal{L_{\text{reg}}},\end{split}$ (16)
@@
4 Experiments
4.1 Experimental Settings
Implementation details: We implement model training and inference on the NVIDIA A800-80GB platform. Specific parameter settings are as follows: the number of training epochs $E$ is set to 500; the input image size $S$ is 128×128; the block numbers of the DRWKV structure are set to $N_{1}=4$ and $N_{2}=8$ respectively, and the input channel number $C$ is 16. The model optimizer uses Adam [kingma2014adam] with parameters set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$; the initial learning rate is $2\times 10^{-4}$, which is gradually reduced to $1\times 10^{-6}$ via the cosine annealing strategy [loshchilov2016sgdr].
Datasets: (1) LSRW-Huawei is a subset of the LSRW [HAI2023103712] dataset consisting of images captured by Huawei smartphones, which includes 2450 pairs of training images and 30 pairs of test images. (2) LOLv2 [9328179] is divided into two subsets: LOLv2 Real and LOLv2 Synthetic. The former is a dataset of real low-light scenes, containing 689 pairs of training images and 100 pairs of test images, covering various environments such as campuses and clubs. The latter is a synthetic dataset, consisting of 900 pairs of training images and 100 pairs of test images. (3) SDSD [9710730] dataset has two subsets: indoor and outdoor. The indoor subset includes 62 pairs of low-light/normal-light videos for training and 6 pairs of low-light/normal-light videos for testing. The outdoor subset contains 116 pairs of low-light/normal-light videos for training and 10 pairs of low-light/normal-light videos for testing.
Metrics: We use traditional quality (distortion) metrics PSNR and SSIM [1284395] to evaluate the results of LLIE, aiming to measure the pixel-level similarity and distortion between the enhanced images and reference images. Additionally, NIQE [6353522] is employed to assess the no-reference perceptual quality of the enhanced images.
4.2 Comparative Experiments
@@
Based on the quantitative results across five test benchmarks, the DRWKV model achieves performance improvements over multiple baseline methods with extremely low GFLOPS, as shown in Table 1. Specifically, DRWKV exhibits superior performance in terms of naturalness on LSRW–Huawei, characterized by low baseline values of NIQE and a high optimization rate (achieves a 1.1% optimization compared to the second-place method). Additionally, on LOLv2, DRWKV shows differences in optimization strengths between the Real and Synthetic test benchmarks: it achieves the optimal model performance with an average improvement of 5.00 dB in PSNR on Real and 0.152 in SSIM on Synthetic, balancing the enhancement of image quality and the preservation of structural information. In low-light scenario videos, by independently enhancing low-light frames in SDSD, the model demonstrates average PSNR differences of up to 9.42 dB/8 dB in indoor and outdoor scenes, ensuring high “restoration accuracy” of image details. For visualization results, refer to Figure 3; the generated results of our DRWKV model exhibit continuous object edges and high-fidelity details, which are highly consistent with the subject-focused edges.
Table 2: Comparison of different Block combinations on the LOLv2 Real benchmark dataset.
@@
Retinex Configuration Results: This experiment focuses on the contradiction of edge distortion under low light. In terms of network structure, replaceable structures Block1 and Block2 are preset to assist in configuring Retinex, enabling DRWKV to achieve global optimum. Since the input end of the GER theory needs to receive the edge feature $E$ (a parameter from the output end of DRWKV), Block1 is constrained by structural limitations. To minimize disturbance to the network architecture, one-sided fixation is ultimately adopted. Deploying the GER theory into Block2 results in remarkable improvement rates of 6.9%/12.3%/4.8% for PSNR, SSIM, and NIQE, respectively, as shown in Table 2. This strongly validates the core idea of gradient-guided global-edge coupling in the GER theory, demonstrating its strong practicality in addressing edge issues of low-light images and further confirming Hypothesis(1).
4.3 Ablation Studies
Loss Performance Results: A targeted ablation study was performed on the proposed loss function combinations. In Figure 4(a), the absence of the decompositional consistency loss $\mathcal{L_{\text{recon}}}$ results in an overall darker reconstructed image with incomplete edge information, indirectly validating its role in structural and luminance reconstruction. In Figure 4(b), removing the edge sparsity loss $\mathcal{L_{\text{sparse}}}$ leads to false edges generated by Gaussian noise contaminating flat regions, highlighting its contribution to preserving and enhancing low-light edges. Figures 4(c) and 4(d) illustrate uneven illumination and over-denoising effects caused by the removal of the illumination smoothness loss $\mathcal{L_{\text{smooth}}}$ and artifact suppression loss $\mathcal{L_{\text{artifact}}}$, respectively, confirming their balancing roles. Figure 4(e) demonstrates model instability with negative image artifacts when the parameter regularization loss $\mathcal{L_{\text{reg}}}$ is removed, indicating that effective regularization can improve the model’s adaptability to unseen data. Additionally, the loss parameters for this model are specified as follows: $\lambda_{1}=1$, $\lambda_{2}=0.01$, $\lambda_{3}=0.1$, $\lambda_{4}=0.05$, and $\lambda_{5}=1e-4$.
@@ 4.3 Ablation Studies
Component Ablation Results: This experiment aims to verify the effectiveness of the Evolving Scanning mechanism proposed in Hypothesis(2) for the edges of low-light images. Structurally, a progressive component integration strategy is adopted, with LOLv2 Real as the test benchmark and the basic structure as the test framework, to construct a five-level progressive validation system. The component configurations, ablation variables, and validation results at each level are shown in Table 3. In terms of the basic structure, we retain the ViT-style block stacking structure of VRWKV and introduce the SSIM+L1 loss, enabling it to initially possess the capability to handle low-light issues. However, the model exhibits nearly ineffective feature extraction in illumination-degraded scenarios, characterized by extremely low SSIM (0.415) and PSNR (12.57 dB) values.
@@
Validation (3): Adding a light preprocessing(LP) stage leads to a slight increase in image quality evaluation metrics, demonstrating that this stage contributes to improving pixel errors and structures.
@@
Validation (5): Replacing the SSIM+L1 loss with the proposed MS2-Loss specifically optimizes the training process, with SSIM and PSNR optimized to 0.832 and 24.12 dB, respectively.
@@
Table 3: Ablation study on the LOLv2 Real benchmark dataset. If the RWKV blocks and training losses are not selected, Bi-RWKV and SSIM+L1 loss will be used by default. ✔ denotes the use of a ViT-style block stacking structure, whereas ✔ indicates the adoption of a UNet-like architecture.
ES-RWKV LP Bi-SAB MS2-Loss SSIM↑ PSNR↑ Params.(M) GFLOPs
@@
4.4 Extended Validation
To evaluate generalization, we applied DRWKV to low-light multi-object tracking on the UAVDark135 benchmark [li2021allday], comprising 135 sequences across diverse scenarios at 30 FPS and 1920×1080 resolution.
@@ 4.4 Extended Validation
Table 4: Extended validation results of DRWKV on low-light object tracking performance. MOTA [bernardin2008evaluating], IDF1 [ristani2016performance] and HOTA [luiten2021hota] are common performance evaluation metrics for object tracking.
@@
ByteTrack [zhang2022bytetrack] 23.434 62.776 50.117 28.623 65.852 52.822
BoT-SORT [aharon2022bot] 18.242 62.916 49.556 25.496 64.049 51.435
Hybrid-SORT [yang2024hybrid] 20.684 62.054 49.196 29.355 66.365 52.905
OC-SORT [Cao_2023_CVPR] 21.062 62.967 49.469 28.102 66.000 52.501
Deep OC-SORT [maggiolino2023deep] 19.514 63.205 49.857 24.929 58.016 49.124
@@
5 Conclusion
In this study, we introduced the DRWKV model to address the challenge of preserving edge details in Low-Light Image Enhancement (LLIE). By integrating the Global Edge Retinex (GER) theory, the model decouples illumination and edge structures, ensuring high-fidelity edge continuity even in extreme low-light conditions. The Evolving WKV Attention mechanism enhances spatial edge continuity, while the Bilateral Spectrum Aligner Block (Bi-SAB) and MS2-Loss optimize luminance and chrominance alignment, reducing artifacts and improving visual quality. Extensive experiments on benchmark datasets show that DRWKV outperforms existing methods in PSNR, SSIM, and NIQE, with lower computational cost. Additionally, it proves effective in tasks like low-light multi-object tracking, enhancing performance, especially in identity preservation. Future work will focus on adapting DRWKV for mobile platforms and further improving its robustness across various low-light scenarios, bridging the gap between academic research and real-world deployment.
DRWKV: Focusing on Object Edges for Low-Light Image Enhancement
@@ 6 Supplementary Experiments
7 Architecture Details of DRWKV
@@ 7.1 Light Preprocessing
Figure 5: Evolving Scanning RWKV’s structure, it consists of two parts: Evolving Scanning Spatial Mix and Channel Mix.
7.2 Evolving Scanning RWKV
Unlike traditional spatial modelling approaches, the proposed ES-RWKV breaks through spatial-size constraints, preserves structural integrity, enhances modelling capacity, and enables cross-space information interaction while markedly reducing the computational overhead of handling complex spatial features, thereby offering an efficient solution for spatial feature extraction and fusion in vision tasks. The ES-RWKV block is decomposed into two sequential stages: Evolving Scanning Spatial Mix and Channel Mix. The Evolving Scanning RWKV structure is shown in Figure.5
7.2.1 Evolving Scanning Spatial Mix